# LIFT: Improving Long Context Understanding of Large Language Models through Long Input Fine-Tuning

### v1

@@ 4.1 Setup
The evaluation metrics are consistent with those used in the original benchmarks. For LongBench and NIAH, the evaluation metrics are task-specific (Zhang et al., 2020). Since most automatic evaluation metrics are sensitive to semantic expression, output format, and length, we utilize GPT4-0613 (Achiam et al., 2023) for LooGLE as recommended in the paper to judge whether the two answers are semantically the same or not, noted as GPT4_score. It has been proven to exhibit high consistency with human evaluation and can serve as a reliable annotator to a great extent (Suri et al., 2023; Liu et al., 2023; Zheng et al., 2023). The prompts implemented can be found in Section A.3.
Models. For open-source LLMs, we select Llama-3-8B-Instruct (Dubey et al., 2024) and Gemma-2-9B (Team, 2024) with 8k context windows. For closed-source commercial LLMs, we choose GPT-3.5-turbo-0125 (Chen et al., 2023a) with a 16k context window. It has shown competitive performance on popular long context benchmarks and can be further fine-tuned through API service. Details of the models and hyper-parameters can be seen in Appendix A.
Settings. We have highlighted the differences between LIFT and existing long-context methods (Tworkowski et al., 2024; Chen et al., 2023b; Yang, 2023) in Table 1. The goal of LIFT is to augment any short-context LLMs with long-context abilities. Therefore, we evaluate whether LIFT can improve over the pure ICL performance of the above LLMs. Specifically: ICL denotes truncating the long input by retaining only the initial and final tokens in the context window and using pure ICL over the truncated context with the original model. LIFT denotes first fine-tuning the LLM using the Gated Memory adapter (except for GPT-3.5 which uses the default API tuning service) with the objectives in Equations (4) and (5), where the number of auxiliary tasks used and the method for generating them are specialized for each subtask (detailed in Appendix A), and then using truncated ICL (also retraining only the initial and final tokens) with the fine-tuned model.
@@ 4.2.1 Results on LooGLE
Overall performance. LooGLE is a very challenging long-context benchmark with human-proposed/verified long-dependency QAs (LongQA) whose answers require retrieving evidences from multiple pieces spanning the whole long input, as well as automatically generated short-dependency QAs (ShortQA). As shown in Table 2, LIFT consistently outperforms ICL often by large margins on both LongQA and ShortQA tasks across the three LLMs. Particularly, it shows notable improvement in GPT4_score from 15.44 to 29.97 on Llama 3 in LongQA and from 37.37 to 50.33 on Gemma 2 in ShortQA. The results highlight that LIFT significantly improves the performance of ICL, particularly for models with short context windows. Notably, GPT-3.5 generally outperforms the open-sourced models across both tasks, while LIFT can further boost its performance. It can be noticed that all the models perform poorly on LongQA, with GPT4_score falling below 50. This underscores that modeling long dependencies in extended contexts remains a significant challenge for existing methods.
@@ 4.2.2 Results on LongBench
Table 3 presents the results across five representative tasks with extremely long inputs in LongBench. We follow the evaluation metrics introduced in the original benchmark for comparison. For Llama 3, LIFT outperforms ICL on 4 out of 5 subtasks, and for GPT-3.5, LIFT outperforms ICL on 3 out 5 subtasks. We make in-depth analysis to figure out the impact of LIFT on different subtasks. NarrativeQA and QMSum consistently have performance gains from LIFT since these two tasks are similar to axiliary tasks in LIFT. In contrast, for tasks that mainly depend on the memorization of long inputs like Musique and PassageRetrievalEN, LIFT’s advantage is not consistent. Empirically, we found it hard to hit a perfect balance between the long input memorization in LIFT and the ICL ability of the original model. When the model is significantly fine-tuned to memorize the long input, its original capabilities tend to degrade more. As most tasks do not require perfect memorization of the long input, our current strategy is to early-stop the fine-tuning to avoid overfitting. As a consequence, Llama 3 struggles to memorize the details in the context, leading to poor performance on Musique, while GPT-3.5 slightly benefits from LIFT, likely due to more robust foundation capabilities and better fine-tuning strategies. On PassageRetrievalEN, since we imitate the test set generation process to synthesize auxiliary tasks, Llama 3 benefits from the auxiliary tasks, becoming more familiar with the instructions at test time. In contrast, GPT-3.5, due to its strong instruction-following capability, does not see significant improvements from LIFT.
@@ 4.2.3 Efficiency
Benefiting from our segmentation strategy (Section 3.1), the computational complexity of our method scales linearly with the input context length. To further evaluate the efficiency of our approach compared to ICL (with full input in the context), we measure the time cost of a single Needle-In-A-Haystack (NIAH) task under both methods. In this experiment, the input lengths are controllable and the primary computational cost stems from processing the input context rather than iterative generation. We compare LIFT with the Gated Memory adapter and another adapter PiSSA (Meng et al., 2024) against ICL, plotting GPU time versus input length along with the fitted curves in Figure 4.
First, we observe that LIFT is significantly more memory-efficient than ICL. Notably, ICL runs out of memory when the input length exceeds 90k tokens on our A800 (80G) system. For ICL, the cache of hidden states of previous tokens consumes most of the memory. In contrast, LIFT is capable of handling arbitrarily long inputs. Our segmentation and truncation strategy ensures that LIFT only involves training and inference with short text segments, eliminating the need for extensive caching.
@@ 4.3 Ablation Study
Contextualized training. As discussed in Section 3.3, we posit that contextualized training is a pivotal component of the LIFT framework. By aligning segmented language modeling and auxiliary tasks within the same semantic space and task format, contextualized training enables the model to memorize input while minimizing the degradation of its inherent capabilities. As demonstrated in Table 4, the inclusion of contextualized training significantly enhances model performance on both the LooGLE ShortQA and LongQA tasks compared to the version without this component. This improvement underscores the critical role of contextualized training in achieving robust and effective long-text understanding.
@@
Datasets PiSSA Gated-Memory
@@ 4.4 Details and Analysis
Table 7: In-context and out-context scores on LooGLE ShortQA. LIFT (PiSSA) denotes the fine-tuning approach using PiSSA instead of the Gated-Memory architecture.
@@
As delineated in Table 7, we have evaluated the GPT_score of the LIFT approach separately for in-context and out-of-context questions. After LIFT, the model’s ICL abilities are inevitably compromised; however, there is a corresponding enhancement in the out-of-context capabilities based on memory. This trade-off inherently limits the efficacy of the naive approach of directly fine-tuning on the long input. It is evident that our LIFT method, by contrast–—when compared to approaches that neither employ contextualized training nor utilize Gated Memory—–significantly mitigates the decline in in-context scores while bolstering the improvement in out-of-context scores. This observation aligns with the motivation outlined in Section 3.3 and underscores the advantages of Gated Memory in better balancing the original ICL ability and the new ability adapted for the long input.
@@ 6 Limitations and Future Work
Strategy to extract parametric knowledge after LIFT. Through LIFT, embedding inputs into the model’s internal parameters enhances its familiarity with the data. However, the effectiveness of downstream tasks still relies on the model’s ability to autonomously extract and utilize the parametric knowledge acquired during LIFT. The detailed analysis (Section 4.4) reveals a significant performance gap between in-context and out-of-context questions, suggesting that the model’s capability to extract parametric knowledge post-LIFT needs further improvement. This presents a promising direction for future research and exploration.
Challenges using LIFT with auxiliary tasks. Our findings reveal that auxiliary tasks during LIFT offer minimal benefit and can even degrade performance due to overfitting. Additionally, simply fine-tuning the model on long texts does not inherently endow it with robust reasoning capabilities over such texts. These observations underscore the necessity for more effective strategies to harness the in-parameter knowledge of LLMs, enabling them to reason efficiently and accurately over extended contexts.

### v2

@@ 4.1 Setup
The evaluation metrics are consistent with those used in the original benchmarks. For LongBench and NIAH, the evaluation metrics are task-specific (Zhang et al., 2020). Since most automatic evaluation metrics are sensitive to semantic expression, output format, and length, we utilize GPT4-0613 (Achiam et al., 2023) for LooGLE as recommended in the paper to judge whether the two answers are semantically the same or not, noted as GPT4_score. It has been proven to exhibit high consistency with human evaluation and can serve as a reliable annotator to a great extent (Suri et al., 2023; Liu et al., 2023; Zheng et al., 2023). The prompts implemented can be found in Appendix A.
Models. For open-source LLMs, we select Llama-3-8B-Instruct (Dubey et al., 2024) and Gemma-2-9B-it (Team, 2024) with 8k context windows. For closed-source commercial LLMs, we choose GPT-3.5-turbo-0125 (Chen et al., 2023a) with a 16k context window. It has shown competitive performance on popular long context benchmarks and can be further fine-tuned through API service. Details of the models and hyper-parameters can be seen in Appendix A.
Settings. In this paper, we mainly evaluate LIFT compared with the truncated ICL performance of the selected LLMs. Specifically: ICL denotes truncating the long input by retaining only the beginning and end of texts within the context window of the original model. LIFT denotes first fine-tuning the LLM using the Gated Memory adapter (except for GPT-3.5 which uses the default API tuning service) with the objectives in Equations (4) and (5) with the above-mentioned truncated ICL. The number of auxiliary tasks used and the method for generating them are specialized for each subtask (detailed in Appendix A).
@@ 4.2.1 Results on LooGLE
Overall performance. As shown in Table 2, LIFT consistently outperforms ICL often by large margins on the overall scores on both LongQA and ShortQA tasks across the three LLMs. Particularly, it shows notable improvement in GPT4_score from 15.44 to 29.97 on Llama 3 in LongQA and from 37.37 to 50.33 on Gemma 2 in ShortQA. The results highlight that LIFT significantly improves the performance of ICL, particularly for models with short context windows. Notably, GPT-3.5 generally outperforms the open-sourced models across both tasks, while LIFT can further boost its performance. It can be noticed that all the models perform poorly on LongQA, with GPT4_score falling below 50. This underscores that modeling long dependencies in extended contexts remains a significant challenge for existing methods.
@@ 4.2.2 Results on LongBench
Table 3 presents the results across five representative tasks with extremely long inputs in LongBench. We follow the evaluation metrics introduced in the original benchmark for comparison. For Llama 3, LIFT outperforms ICL on 4 out of 5 subtasks, and for GPT-3.5, LIFT outperforms ICL on 3 out 5 subtasks. We make in-depth analysis to figure out the impact of LIFT on different subtasks. NarrativeQA and QMSum consistently have performance gains from LIFT since these two tasks are similar to auxiliary tasks in LIFT. In contrast, for tasks that mainly depend on the memorization of long inputs like Musique and PassageRetrievalEN, LIFT’s advantage is not consistent. Empirically, we found it hard to hit a perfect balance between the long input memorization in LIFT and the ICL ability of the original model. When the model is significantly fine-tuned to memorize the long input, its original capabilities tend to degrade more. As most tasks do not require perfect memorization of the long input, our current strategy is to early-stop the fine-tuning to avoid overfitting. As a consequence, Llama 3 struggles to memorize the details in the context, leading to poor performance on Musique, while GPT-3.5 slightly benefits from LIFT, likely due to more robust foundation capabilities and better fine-tuning strategies. On PassageRetrievalEN, since we imitate the test set generation process to synthesize auxiliary tasks, Llama 3 benefits from the auxiliary tasks, becoming more familiar with the instructions at test time. In contrast, GPT-3.5, due to its strong instruction-following capability, does not see significant improvements from LIFT.
@@ 4.2.3 Efficiency
Benefiting from our segmentation strategy (Section 3.1), the computational complexity of our method is linear with respect to the input length. To further evaluate the efficiency of our approach compared to ICL (with full input in the context), we measure the time cost of a single Needle-In-A-Haystack (NIAH) task under both methods. In this experiment, the input lengths are controllable and the primary computational cost stems from processing the input context rather than iterative generation. We compare LIFT with the Gated Memory adapter and another adapter PiSSA (Meng et al., 2024) against ICL, plotting GPU time versus input length along with the fitted curves in Figure 4.
First, we observe that LIFT is significantly more memory-efficient than ICL. Notably, ICL runs out of memory when the input length exceeds 90k tokens on our A800 (80G) system. For ICL, the KV cache consumes most of the memory. In contrast, LIFT is capable of handling arbitrarily long inputs. Our segmentation and truncation strategy ensures that LIFT only involves training and inference with short text segments, eliminating the need for extensive caching.
@@ 4.3 Ablation Study
Contextualized training. As discussed in Section 3.3, we posit that contextualized training is a pivotal component of the LIFT framework. By aligning input memorization tasks and auxiliary tasks within the same semantic space and task format, contextualized training enables the model to memorize input while minimizing the degradation of its inherent capabilities. As demonstrated in Table 4, the inclusion of contextualized training significantly enhances model performance on both the LooGLE ShortQA and LongQA tasks compared to the version without this component. This improvement underscores the critical role of contextualized training in achieving robust and effective long-text understanding.
@@
Datasets PiSSA Gated Memory
@@ 4.4 Details and Analysis
Table 7: In-context and out-context scores on LooGLE ShortQA. LIFT (PiSSA) denotes the fine-tuning approach using PiSSA instead of the Gated Memory architecture.
@@
We evaluated the GPT4_score of the LIFT approach separately for in-context and out-of-context questions noted as in- and out-context score respectively in Table 7. After LIFT, the model’s ICL abilities are inevitably compromised; however, there is a corresponding enhancement in the out-of-context capabilities based on memory. This trade-off inherently limits the efficacy of the naive approach of directly fine-tuning on the long input. It is evident that our LIFT method, by contrast–—when compared to approaches that neither employ contextualized training nor utilize Gated Memory—–significantly mitigates the decline in in-context scores while bolstering the improvement in out-of-context scores. This observation aligns with the motivation outlined in Section 3.3 and underscores the advantages of Gated Memory in better balancing the original ICL ability and the new ability adapted for the long input.
@@ 6 Limitations and Future Work
Strategy to extract parametric knowledge after LIFT. Through LIFT, embedding inputs into a model’s internal parameters enhances its familiarity with the data. However, the effectiveness on downstream tasks still relies on the model’s ability to autonomously extract and utilize the parametric knowledge acquired during LIFT. The detailed analysis (Section 4.4) reveals a significant performance gap between in-context and out-of-context questions, suggesting that the model’s capability to extract parametric knowledge post-LIFT needs further improvement. This presents a promising direction for future research and exploration.
Challenges using LIFT with auxiliary tasks. Our findings reveal that auxiliary tasks are beneficial for improving LLMs’ understanding of inputs. However, the benefit comes with extensive computational costs and they must be aligned with test tasks to achieve optimal performance. Future research should explore general auxiliary tasks that are suitable for all types of long inputs and tasks and are more computationally efficient to be incorporated into LIFT.
@@ A.3 Generating auxiliary tasks
A.4 GPT4_score evaluation
We utilize GPT-3.5-turbo-0125 to evaluate the correctness of the responses of LLMs based on the questions and the ground-truth answers on LooGLE.
The prompt is as following (in the format of the chat template):
System: Given one question, there is a groundtruth and a predict_answer. Please decide whether they are the same or not in semantic. Please only output ’True’ or ’False’ . User: Question: {Question} groundtruth = {Ground-truth answer} predict_answer = {LLM response}